1 Keyer; what is a keyer...

Uniambic, biambic...

2 Custom made keyer

A plastic material of relatively low melting temperature is heated in boiling water, and applied to hand, to form to shape of hand. See Figure 1.

Next the soft heated plastic is shaped to perfectly fit the hand. See Fig 2.

Finally, excess material is trimmed away. See Fig 3.

2.1

See Fig 4.

(a)  (b)  (c)

(d)  (e)  (f)

(g)  (h)  (i)

(j)  (k)  (l)

Figure 1: Preparation for making of keyer: (a) pouring water from kettle; (b) putting low temp plastic into hot water; (c) wooden chopsticks are used to handle hot plastic; (d) soft plastic is tested for consistency; (e) hand is cooled in icewater; (f-k) soft plastic is lifted out of hot water in one quick movement while hand is taken out of ice bath, (l) finishing up with soft plastic on hand. If pain threshold reached, process can be repated with further cooling of hand.

(a)  (b)

(c)  (d)

(e)  (f)

(g)  (h)

Figure 2: Shaping the soft heated plastic to perfectly fit the hand: (a) here a left handed keyer is in the making; (b-c) soft heated plastic is pressed down to form nicely between the fingers, so that the keyer will conform perfectly to hand shape. (d-h) when right hand taken away, material should feel comfortable and naturally stay in position on left hand.

(a)  (b)  (c)

(d)  (e)  (f)

(g)  (h)  (i)

(j)  (k)  (l)

Figure 3: Trimming the keyer support structure; when the plastic cools and hardens, it will fit perfectly to the hand, in the interior region of the plastic. However, at the edges, the shape will be whatever was set forth by the original shape of the material, and will therefore need to be trimmed: (a-h) initial rough trimming is done while the support structure is worn, so that it can be trimmed to the hand shape; (i-l) then it is removed and rough edges are rounded of nicely.

(a)  (b)

(c)  (d)

(e)  (f)

(g)  (h)

Figure 4: Finishing off a keyer: (a) selecting appropriate pushbutton switches; (b) testing for positioning of the switches; (c) getting the thumb switches in position is most important; (d) material is cut away and switches should be glued into place all while wearing the keyer; (e) keyer is interfaced to WearComp by way of a Programmable Interface Controller, using a circuit board designed by Wozniak (assembled by author) and a PIC program designed by Wozniak, Mann, Moncrieff, and Fung; (f) circuit board (in author’s right hand) interfaces key switches to standard PS/2 connector (in author’s left hand); (g) an RJ45 connector is used with 8 wires, one for each of the seven switches and the eighth for common; (h) finally the keyer can be plugged in by way of the RJ45 connector. The author has standardized the connections so that any member of the community having difficulty with their keyer can temporarily swap with someone else to determine whether the problem is in the keyer or in the conversion circuit.